if you invested $250 at 16% interest how much will you have after 18 years?
Answer by Alan3354(69443)

if you invested $250 at 16% interest how much will you have after 18 years?
-----------------
If it's simple interest, you'll have $250*0.16*18 = $720
------------
If it's compount interest, which is likely, it varies with the compounding period.
---
Compounded annually it's $250*(1.16)^18 = $3615.63
Compounded quarterly it's $250*(1.04)^72 = $4210.57
Compounded continuously it's $250*e^0.16*18 = $ 4453.68